find the vertex of the parabola: y=4x^2-8x+5
Answer by khwang(438)

Make the complete square
y=4x^2-8x+5 = 4(x^2 -2x +1) -4 + 5
= 4(x-1)^2 + 1
OR y - 1 = 4(x-1)^2
Hence, the vertex of this parabola is (1,2) & cancave upward.